OpenStudy (anonymous):

solve the following: x=6+3z y+2z=2 7x-5z=3y+14

OpenStudy (anonymous):

once you substitute x and y into 7x - 5z = 3y + 14 you get 42 + 21z = 6 -6z + 14

OpenStudy (anonymous):

that's it?

OpenStudy (anonymous):

well, no, you're solving for X, Y, and Z right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

help me to finish it please

OpenStudy (anonymous):

i just realize i left out a -5z so messed everything up lol.

OpenStudy (anonymous):

Z = 1 can you find the rest or you still need help?